On March 17, 2017, Ann and Bob agree to form Robin Corporation. Ann transfers a capital asset worth $420,000 (basis of $150,000) for 70 shares in Robin Corporation, with a value of $420,000. Bob transfers inventory worth $180,000 (basis of $150,000) and receives 30 shares, with a value of $180,000. Ann's holding period for the capital asset began on 07/06/14 and Bob's holding period for the inventory began 01/20/17. What gain or income, if any, will be realized on the transfer by Ann 0 and Bob 15,000 ? What gain or income, if any, will be recognized on the transfer by Ann and ? What basis do Ann 150,000 and Bob 45,000 have in the Robin Corporation stock? What is Robin Corporation’s basis in the property and it received from Ann 150,000 and Bob 30,000 ? When do Ann and Bob's holding period start for the Robin Corporation stock? Enter date as xx/xx/xx. When does Robin Corporations holding period start for the capital asset and for the inventory ? Enter date as xx/xx/xx.
Answer #1

1. Ann and Bob do not recognize any loss on the property transferred since no cash is received for the assets transferred.

2. Ann's basis = 150,000

Bob's basis = 150,000

3. Robin corporation's basis in the property = $150,000 +150,000 = 300,000

4. Ann's and Bob's holding period = 03/17/2017

5. Robin corporations holding period:

Capital asset = 07/06/14

Inventory = 01/20/17

The question is slightly distorted and hence have answered in the best possible way.

Any doubt please comment.